At an ice cream parlor, ice cream cones cost $1.10 and sundaes cost $2.35. One day the receipts for a total of 172 cones and sundaes were $294.20. How many cones were sold?
Answer by checkley71(8403)

C+S=172 & 1.10C+2.35S=294.20
C=172-S SUBSTITUTING THIS VALUE OF C IN THE SECOND FORMULA THEN WE GET
1.10(172-S)+2.35S=294.20 OR 189.20-1.10S+2.35S=294.20 OR
-1.10S+2.35S=294.20-189.20 OR 1.25S=105 OR S=105/1.25 OR S=84 THEREFORE
84+C=172 OR C=172-84 OR C=88.
PROOF 84 TIMES 2.35 = $197.40
88 TIMES 1.10 = $ 96.80
-------
$294.20